OpenStudy (juan1857):

find the critical points of y=(x^2)/(4x+4)

OpenStudy (danjs):

Do you know what you have to do?

OpenStudy (juan1857):

take the derivative, and set it equal to 0

OpenStudy (user123):

Yup.

OpenStudy (user123):

Can you find the derivative of that?

OpenStudy (user123):

u'v - uv' / v^2 Remember the denominator is u and the numerator is v.

OpenStudy (juan1857):

yes, i believe it is (4x^2+8)/(4x+4)^2

OpenStudy (danjs):

looks almost right... \[\large y'=\frac{ (4x+4)(2x)-x^2(4) }{ (4x+4)^2 }=\frac{ 4x^2+8x }{ (4x+4)^2 }\]

OpenStudy (danjs):

\[\large = \frac{ 4x(x+2) }{ 16(x+1)^2 }=\frac{ x(x+2) }{ 4(x+1)^2 }\]

OpenStudy (juan1857):

so the points would be -2,-1, and 0??

OpenStudy (danjs):

right, the x values to test

OpenStudy (juan1857):

ok thanks :)

OpenStudy (danjs):

test to see which ones if any are crit points

OpenStudy (juan1857):

oh an um do we seperatley equal them to 0? for example x(x+2)=0 or 4(x+1)^2 or do we do it as a whole?

OpenStudy (danjs):

A Critical Point at x=c says, f(c) exists f '(c)=0 or f '(c) doesnt exist

OpenStudy (danjs):

The original function has x=-1 not existing. All other values are fine for x. So only the two points (-2 , -1) and (0,0) are critical points